# Traveling supersolid stripe patterns in spin-orbit-coupled Bose-Einstein condensates

### v1

@@ IV.3 Results for the observables
with $N_{v_{x}}$ defined in Eq. (68). Both the energy per particle and the chemical potential are even functions of $s_{p}$, reaching their minimum at $s_{p}=0$, where they decrease to the values derived in Ref. [58]. At small $s_{p}$, they increase quadratically [see Figs. 3(a1)-(a2) and 3(b1)-(b2)]. In contrast, both $\Delta\mu$ and $v_{x}$ are odd functions of $s_{p}$, and grow linearly with positive slope for small $|s_{p}|$, as shown in Figs. 3(c1)-(c2) and 3(d1)-(d2). In the limit $s_{p}\to\pm 1$, the energy per particle approaches that of the plane-wave phase [70, 67]. The behavior of $E/N$ as a function of $s_{p}$ depends strongly on the Raman coupling strength. At low $\Omega_{R}$, the quantity $E/N$ increases monotonically with $|s_{p}|$ and attains its maximum at $s_{p}=\pm 1$, as seen in Fig. 3(a1). In contrast, at large $\Omega_{R}$, the quantity $E/N$ develops a global maximum at an intermediate value of $|s_{p}|$ and decreases thereafter, reaching local minima of the energy at $s_{p}=\pm 1$ corresponding to the plane-wave states [see Fig. 3(a2)]. Due to the thermodynamic relation in Eq. (35), this qualitative change also affects the behavior of $\Delta\mu$ and $v_{x}$. While the chemical potential $\mu$ remains a convex function of $s_{p}$, $\Delta\mu$ at large $\Omega_{R}$ develops two stationary extrema, a minimum at negative $s_{p}$ and a maximum at positive $s_{p}$, corresponding to inflection points of the energy per particle. Both $\Delta\mu$ and $v_{x}$ vanish at the local maxima of $E/N$. Interestingly, in the large-$|s_{p}|$ regime, the fringe velocity may acquire the sign opposite to that of the momentum. However, this regime is dynamically unstable, as discussed in Sec. V.2. Finally, within the metastable window between the critical point $\Omega_{\mathrm{cr1}}$ and the spinodal point $\Omega_{\mathrm{sp}}$ the stationary stripe configuration at $s_{p}=0$ represents only a local minimum of the energy per particle, while the plane-wave states at $s_{p}=\pm 1$ are energetically favored. Note that, although $\mu$ and $\Delta\mu$ become ill defined individually at $s_{p}=\pm 1$, the combinations $\mu+\Delta\mu/2$ (for $s_{p}\to 1$) and $\mu-\Delta\mu/2$ (for $s_{p}\to-1$) smoothly approach the chemical potential of the plane-wave phase [67].
@@ V Excitation spectrum of moving stripe patterns
In this section, we examine dynamical properties of moving stripe patterns. We begin by outlining the Bogoliubov theory for studying small oscillations around the stripe background in the comoving frame, where the fringes remain stationary (Sec. V.1). We then compute the Bogoliubov spectrum and the associated sound velocities, highlighting the emergence of both energetic and dynamical instabilities at large momenta (Sec. V.2).
@@ V.1 Bogoliubov theory in the comoving frame
We study the dynamical properties of moving stripe patterns in the comoving frame. Within the framework of Bogoliubov theory [80, 81, 87] small oscillations around the stationary configuration (12) are described by an order parameter of the form
$\Psi^{\prime}(\bm{r}^{\prime},t^{\prime})=\mathrm{e}^{-\mathrm{i}\mu^{\prime}t^{\prime}/\hbar}\mathrm{e}^{-\mathrm{i}m\bm{v}\cdot\bm{r}^{\prime}/\hbar}\left[\Psi_{0}(\bm{r}^{\prime})+\delta\Psi(\bm{r}^{\prime},t^{\prime})\right]\,.$ (37)
To determine the small fluctuation $\delta\Psi(\bm{r}^{\prime},t^{\prime})$, we insert the Ansatz (37) into the Gross-Pitaevskii equation (10). Retaining only terms linear in $\delta\Psi$, we obtain the evolution equation
$\mathrm{i}\hbar\partial_{t^{\prime}}\delta\Psi=\left(h_{\mathrm{SO}}-\mu-\bm{v}\cdot\bm{p}+h_{\mathrm{D}}\right)\delta\Psi+h_{\mathrm{C}}\delta\Psi^{*}\,,$ (38)
@@
$\begin{split}h_{\mathrm{D}}={}&{}g_{dd}(\Psi_{0}^{\dagger}\Psi_{0}\mathbb{I}_{2}+\Psi_{0}\Psi_{0}^{\dagger})\\ &{}+g_{ss}\left[(\Psi_{0}^{\dagger}\sigma_{z}\Psi_{0})\sigma_{z}+(\sigma_{z}\Psi_{0})(\sigma_{z}\Psi_{0})^{\dagger}\right]\,,\end{split}$ (39a)
$h_{\mathrm{C}}={}$ ${}g_{dd}\Psi_{0}\Psi_{0}^{T}+g_{ss}(\sigma_{z}\Psi_{0})(\sigma_{z}\Psi_{0})^{T}\,.$ (39b)
As is customary, we look for solutions of the linearized problem (38) with a single, possibly complex, oscillation frequency:
$\delta\Psi(\bm{r}^{\prime},t^{\prime})=U(\bm{r}^{\prime})\mathrm{e}^{-\mathrm{i}\omega t^{\prime}}+V^{*}(\bm{r}^{\prime})\mathrm{e}^{\mathrm{i}\omega^{*}t^{\prime}}\,.$ (40)
The two-component amplitudes $U$ and $V$ depend on position and are defined up to a global complex normalization factor. Substituting Eq. (40) into Eq. (38) and collecting terms oscillating at the same Bogoliubov frequency $\omega$ we obtain an eigenvalue equation for $\omega$ and the corresponding amplitudes:
$(\eta\mathcal{B}-\bm{v}\cdot\bm{p})\begin{pmatrix}U\\ V\end{pmatrix}=\hbar\omega\begin{pmatrix}U\\ V\end{pmatrix}\,,$ (41)
@@
These expressions formally coincide with those derived for the stationary stripe phase [76, 58], except for the additional Doppler term $-\bm{v}\cdot\bm{p}$ on the left hand side of Eq. (41). As in the $\bm{v}=0$ case, we seek solutions in the form of Bloch waves:
$\begin{pmatrix}U_{b,\bm{k}}(\bm{r}^{\prime})\\ V_{b,\bm{k}}(\bm{r}^{\prime})\end{pmatrix}=\frac{\mathrm{e}^{\mathrm{i}\bm{k}\cdot\bm{r}^{\prime}}}{\sqrt{V}}\sum_{\text{{\hbox{\bar{m}}} odd}}\begin{pmatrix}\tilde{U}_{b,\bm{k},\bar{m}}\\ \tilde{V}_{b,\bm{k},\bar{m}}\end{pmatrix}\mathrm{e}^{\mathrm{i}\bar{m}k_{1}x^{\prime}}\,.$ (43)
Here, $\bm{k}$ is the excitation quasimomentum, while the $\tilde{U}_{b,\bm{k},\bar{m}}$’s and $\tilde{V}_{b,\bm{k},\bar{m}}$’s are two-component expansion coefficients. Combining Eqs. (41) and (43) we obtain an infinite-dimensional eigenvalue problem for these coefficients and the corresponding Bogoliubov frequencies $\omega_{b,\bm{k}}$. A detailed discussion of the case $\bm{v}=0$ can be found in Ref. [54]. At fixed $\bm{k}$, one obtains an infinite set of eigenfrequencies that form a band structure as $\bm{k}$ varies [76, 58]. This justifies the use of an additional subscript, the band index $b$, to label the amplitudes and the frequency of each Bogoliubov mode. The most general solution to Eq. (38) can then be expressed as a sum over all physically distinct excitation bands and quasimomenta.
@@
    complex frequencies occur in complex conjugate pairs. This follows from the fact that the operator $\eta\mathcal{B}-\bm{v}\cdot\bm{p}$ differs from its Hermitian adjoint by a unitary transformation: $\eta(\eta\mathcal{B}-\bm{v}\cdot\bm{p})\eta^{-1}=(\eta\mathcal{B}-\bm{v}\cdot\bm{p})^{\dagger}$ [87];
@@
    each mode has a conjugate counterpart that describes the same physical oscillation. Specifically, if the pair $(U_{b,\bm{k}},V_{b,\bm{k}})$ solves Eq. (41) with frequency $\omega_{b,\bm{k}}$, then the pair $(V_{b,\bm{k}}^{*},U_{b,\bm{k}}^{*})$ is a solution with frequency $-\omega_{b,\bm{k}}^{*}$ [87]. However, because of the structure of the Ansatz (40), these two solutions describe the same physical excitation;
@@
$\begin{split}&{}\int_{V}\mathrm{d}^{3}r^{\prime}\,\left[U_{b^{\prime},\bm{k}^{\prime}}^{\dagger}(\bm{r}^{\prime})U_{b,\bm{k}}(\bm{r}^{\prime})-V_{b^{\prime},\bm{k}^{\prime}}^{\dagger}(\bm{r}^{\prime})V_{b,\bm{k}}(\bm{r}^{\prime})\right]\\ &{}=\mathcal{N}_{b,\bm{k}}\delta_{bb^{\prime}}\delta_{\bm{k}\bm{k^{\prime}}}\,,\end{split}$ (44)
which has two key implications. First, it defines orthogonality between modes belonging to different bands ($b^{\prime}\neq b$) or with different quasimomenta ($\bm{k}^{\prime}\neq\bm{k}$). Second, in the case $b^{\prime}=b$ and $\bm{k}^{\prime}=\bm{k}$, it provides the norm $\mathcal{N}_{b,\bm{k}}$ of a given Bogoliubov mode. Importantly, each physical oscillation corresponds to a pair of solutions with opposite norms. In plotting the Bogoliubov spectrum, such as in Fig. 4, it is customary to include only the frequencies of the positive norm solutions. Furthermore, the system is said to be energetically stable if all positive norm solutions have positive frequency. In the next section, we will show that moving stripe patterns can lose energetic or dynamical stability at sufficiently large momentum. Finally, due to the periodicity of the spectrum, it suffices to consider values of the $x$ component of $\bm{k}$ within the first Brillouin zone. In contrast, the $y$ and $z$ components are unconstrained. Following the conventions of Refs. [76, 58], we define the first Brillouin zone for the excitation quasimomentum as the interval $0\leq k_{x}<2k_{1}$. With this choice, excitations propagating along the positive $x$ direction satisfy $0<k_{x}<k_{1}$, while those along the negative $x$ direction correspond to $k_{1}<k_{x}<2k_{1}$.
@@ V.2 Bogoliubov spectrum and sound velocities
To numerically compute the Bogoliubov spectrum, we truncate the expansions (43) by retaining only a finite number of terms. Specifically, we include components with $-9\leq\bar{m}\leq 9$, following the approach used in Sec. IV for evaluating $\Psi_{0}$. This truncation reduces Eq. (41) to a finite-dimensional eigenvalue problem, which can be readily solved. A detailed discussion of the computational procedure is given in Ref. [54]. In Fig. 4, we display the four lowest-lying bands of the computed spectrum for excitation quasimomentum $\bm{k}$ aligned along the $x$ axis. The figure consists of two rows of panels, corresponding to the two values of the Raman coupling considered in this work. Within each row, we vary the value of $s_{p}$ to illustrate four representative cases:
Figure 4: Excitation spectrum of a spin-orbit-coupled BEC in the stripe phase for various values of $s_{p}$: (a1)-(a2) $s_{p}=0$ (stationary stripes), (b1)-(b2) $s_{p}=0.5$, (c1) $s_{p}=0.95$, (c2) $s_{p}=0.75$, and (d1)-(d2) $s_{p}=0.99$. The top-row panels correspond to $\hbar\Omega_{R}/E_{R}=1.0$, while those in the bottom row correspond to $\hbar\Omega_{R}/E_{R}=2.6$. The excitation quasimomentum is taken along the $x$ direction. Only the four lowest-lying excitation bands are shown; the band index $b$ is indicated next to each curve. In (c1), the inset shows a magnified view near the left edge of the Brillouin zone, where the lowest gapless band exhibits a negative frequency. In all panels, we plot only the real part of the frequency, $\omega_{b,k_{x}}^{\mathrm{R}}$; the imaginary part, $\omega_{b,k_{x}}^{\mathrm{I}}$, is zero except for the lowest gapless band in (c2) (see inset) and in (d2), where it is nonzero but very small and omitted for clarity. The red dashed line in (c2) displays the solution $\omega_{1,\bm{k}}^{*}$ in the dynamically unstable region and $-\omega_{1,2k_{1}\hat{\bm{e}}_{x}-\bm{k}}$ in the dynamically stable region, both obtained from Eq. (41). In panels (d1)-(d2), the black dashed curves represent the two branches of the excitation spectrum in the plane-wave phase, computed using the same parameters as in the case of the stripe-phase spectrum, and with an added Doppler shift (see text for details). The interaction parameters are as in previous figures: $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$.
@@
    Spectrum of stationary stripes ($s_{p}=0$), previously computed in Refs. [76, 58]. The two lowest bands are gapless and exhibit linear dispersion near both edges of the first Brillouin zone, i.e., at $k_{x}=0$ and $k_{x}=2k_{1}$. The two zero-frequency modes at $k_{x}=0$ correspond to the Goldstone modes arising from the spontaneous breaking of global-phase and translation symmetries. Recall that, in the regime $g_{dd}>g_{ss}$ considered here, the upper (lower) phonon branch has predominantly density (spin) character [76], corresponding to the superfluid (crystalline) nature of the stripe phase [58]. The roles are interchanged when $g_{ss}>g_{dd}$.
@@
    Spectrum of moving stripes at intermediate momentum ($s_{p}=0.5$). Most features remain similar to the stationary case. However, the bands no longer exhibit reflection symmetry with respect to the center of the Brillouin zone, $k_{x}=k_{1}$. This is due to the fact that the transformation $k_{x}\to 2k_{1}-k_{x}$ transforms the spectrum of a stripe pattern to the one with opposite momentum. Additionally, the gaps between bands are narrower for moving stripes, and this trend becomes more pronounced as $|s_{p}|$ increases.
@@
    Spectrum in the instability regime ($s_{p}=0.95$ for the top row, $s_{p}=0.75$ for the bottom row). The behavior depends on the Raman coupling. At lower Raman coupling (top row), a portion of the lowest gapless band dips into negative frequency, particularly near the left edge of each Brillouin zone (or the right edge for negative $s_{p}$), as shown in the inset of panel (c1). This signals the onset of energetic instability for moving stripes at large momenta. In contrast, at higher Raman coupling (bottom row), the system first exhibits energetic instability (now near the right edge of the Brillouin zone), followed by the emergence of a nonzero imaginary part in the lowest band near both edges of the Brillouin zone, as seen in panel (c2). According to the discussion in the previous section, this indicates that moving stripe patterns become dynamically unstable at high Raman coupling and high momentum.
@@
    Spectrum near the melting limit ($s_{p}=0.99$). Unlike the previous panels, which focus on the first Brillouin zone, here we plot the spectrum over the range $-4k_{1}\leq k_{x}<2k_{1}$, encompassing three consecutive Brillouin zones. One observes a significant narrowing of the band gaps, heralding their complete closure at the stripe melting point. At this point, segments from different bands and Brillouin zones merge to form the dispersion relations $\omega_{\pm,\bm{k}}-\bm{v}\cdot\bm{k}$ (dashed lines), where $\omega_{\pm,\bm{k}}$ denote the lower (--) and upper ($+$) branches of the excitation spectrum in the plane-wave phase [67]. [^3] Figure 4(d1)-(d2) may suggest that the Bogoliubov frequencies remain periodic in $k_{x}$ even in the plane-wave phase. However, one must recall that $\bm{k}$ represents the excitation quasimomentum. The spectrum loses periodicity when plotted as a function of the excitation momentum, which is well defined only for $|s_{p}|=1$; see the related discussion in Ref. [58]. The Doppler shift $-\bm{v}\cdot\bm{k}$ appears because the spectrum is evaluated in the comoving frame. Notably, the density phonon modes (upper gapless branch) of the stripe phase smoothly evolve into the single phonon branch of the plane-wave phase. More intriguingly, the spin phonon modes (lower gapless branch) transform into the well-known roton minimum of the plane-wave excitation spectrum near $k_{x}=-2k_{1}^{\mathrm{PW}}$ (or $k_{x}=+2k_{1}^{\mathrm{PW}}$ for the plane-wave state with negative condensation momentum, corresponding to $s_{p}=-1$[88, 67, 60, 89, 90].

### v2

@@ IV.2 Results for the observables at finite Raman coupling
with $N_{v}$ defined in Eq. (83). Both $\bar{\varepsilon}$ and $\mu_{d}$ are even functions of $s_{p}$, reaching their minimum at $s_{p}=0$, where they decrease to the values derived in Ref. [60]. At small $s_{p}$, they increase quadratically [see Figs. 4(a1)-(a2) and 4(b1)-(b2)]. In contrast, both $\mu_{s}$ and $v$ are odd functions of $s_{p}$, and grow linearly with positive slope for small $|s_{p}|$, as shown in Figs. 4(c1)-(c2) and 4(d1)-(d2). In the limit $s_{p}\to\pm 1$, the average energy density approaches that of the plane-wave phase [72, 69]. The behavior of $\bar{\varepsilon}$ as a function of $s_{p}$ depends strongly on the Raman coupling strength. At low $\Omega_{R}$, the quantity $\bar{\varepsilon}$ increases monotonically with $|s_{p}|$ and attains its maximum at $s_{p}=\pm 1$, as seen in Fig. 4(a1). In contrast, at large $\Omega_{R}$, the quantity $\bar{\varepsilon}$ develops a global maximum at an intermediate value of $|s_{p}|$ and decreases thereafter, reaching local minima at $s_{p}=\pm 1$ corresponding to the plane-wave states [see Fig. 4(a2)]. Due to the thermodynamic relation in Eq. (23b), this qualitative change also affects the behavior of $\mu_{s}$ and $v$. While the mean chemical potential $\mu_{d}$ remains a convex function of $s_{p}$, $\mu_{s}$ at large $\Omega_{R}$ develops two stationary extrema, a minimum at negative $s_{p}$ and a maximum at positive $s_{p}$, corresponding to inflection points of the average energy density. Both $\mu_{s}$ and $v$ vanish at the local maxima of $\bar{\varepsilon}$. Interestingly, in the large-$|s_{p}|$ regime, the fringe velocity may acquire the sign opposite to that of the dressed spin imbalance. However, this regime is dynamically unstable, as discussed in Sec. V.2. Finally, within the metastable window between the critical point $\Omega_{\mathrm{cr1}}$ and the spinodal point $\Omega_{\mathrm{sp}}$ the stationary stripe configuration at $s_{p}=0$ represents only a local minimum of the energy per particle, while the plane-wave states at $s_{p}=\pm 1$ are energetically favored. Note that, although $\mu_{d}$ and $\mu_{s}$ become ill defined individually at $s_{p}=\pm 1$, the combinations $\mu_{d}+\mu_{s}$ (for $s_{p}\to 1$) and $\mu_{d}-\mu_{s}$ (for $s_{p}\to-1$) smoothly approach the chemical potential of the plane-wave phase [69].
@@ V Excitation spectrum of moving stripe patterns
In this section, we examine dynamical properties of moving stripe patterns. We begin by outlining the Bogoliubov theory for studying small oscillations around the stripe background in the comoving frame, where the fringes remain stationary (Sec. V.1). We then compute the Bogoliubov spectrum and the associated sound velocities, highlighting the emergence of both energetic and dynamical instabilities at large dressed spin imbalance (Sec. V.2).
@@ V.1 Bogoliubov theory in the comoving frame
The analysis in the previous sections was carried out in the laboratory frame, that is, the rest frame of the Raman lasers responsible for generating the spin-orbit coupling. An equivalent description can be formulated in the comoving frame, which corresponds to the rest frame of the crystal pattern. The space and time coordinates in the two frames are related by the usual transformations $x^{\prime}=x-vt$, $y^{\prime}=y$, $z^{\prime}=z$, and $t^{\prime}=t$ (throughout this work, primed quantities refer to the comoving frame). We have already seen that the order parameter of a moving stripe pattern becomes stationary when expressed in terms of $\bm{r}^{\prime}=(x^{\prime},y^{\prime},z^{\prime})$ and $t^{\prime}$. Consequently, the comoving frame provides the most suitable framework in which to study the dynamical properties of traveling stripes, as their excitation spectrum is well defined in this frame.
$h_{\mathrm{SO}}^{\prime}=h_{\mathrm{SO}}-v\hbar k_{R}\sigma_{z}\,.$ (46)
This extra term can be understood by noting that in the comoving frame the Raman lasers move with velocity v-v along $x$, see Fig. 1(b2). As a result, the frequency of the light field experienced by the atoms undergoes a Doppler shift of $2vk_{R}$, which leads to the corresponding shift in the Raman detuning $\delta_{R}$ by the same amount [62, 86].
In order to apply the Bogoliubov theory [83, 84, 87] and study small oscillations on top of a moving stripe pattern, we write the order parameter in the comoving frame in the form
$\Psi^{\prime}(\bm{r}^{\prime},t^{\prime})=\mathrm{e}^{-\mathrm{i}\mu^{\prime}t^{\prime}/\hbar}\mathrm{e}^{-\mathrm{i}mvx^{\prime}/\hbar}\left[\Psi_{0}(x^{\prime})+\delta\Psi(\bm{r}^{\prime},t^{\prime})\right]\,$, (47)
with $\mu^{\prime}=\mu+mv^{2}/2$. Substituting this expression into Eq. (45), and neglecting for a moment the small fluctuation term $\delta\Psi$, one recovers the time-independent equation (18) obeyed by $\Psi_{0}$. This confirms the consistency of the descriptions in the laboratory and comoving frames. The presence of the $\mathrm{e}^{-\mathrm{i}mvx^{\prime}/\hbar}$ factor in Eq. (47) indicates that in the comoving frame superflow occurs with velocity v-v along the $x$ direction, as shown in Fig. 1(b2).
The linear terms in $\delta\Psi$ that emerge after inserting the Ansatz (47) into the Gross-Pitaevskii equation (45) yield an evolution equation for the small fluctuation,
$\mathrm{i}\hbar\partial_{t^{\prime}}\delta\Psi=\left(h_{\mathrm{SO}}-\mu-vp_{x}+h_{\mathrm{D}}\right)\delta\Psi+h_{\mathrm{C}}\delta\Psi^{*}\,$, (48)
@@
[…]
$h_{\mathrm{C}}={}$ ${}g_{dd}\Psi_{0}\Psi_{0}^{T}+g_{ss}(\sigma_{z}\Psi_{0})(\sigma_{z}\Psi_{0})^{T}\,$, (49b)
and $\mathbb{I}_{2}$ is the $2\times 2$ identity matrix. As is customary, we look for solutions of the linearized problem (48) with a single, possibly complex, oscillation frequency:
$\delta\Psi(\bm{r}^{\prime},t^{\prime})=U_{b,\bm{k}}(\bm{r}^{\prime})\mathrm{e}^{-\mathrm{i}\omega_{b,\bm{k}}t^{\prime}}+V_{b,\bm{k}}^{*}(\bm{r}^{\prime})\mathrm{e}^{\mathrm{i}\omega_{b,\bm{k}}^{*}t^{\prime}}\,.$ (50)
The two-component amplitudes $U_{b,\bm{k}}$ and $V_{b,\bm{k}}$ depend on position and are defined up to a global complex normalization factor. The meaning of the indices $b$ and $\bm{k}$ will be clarified shortly. Substituting Eq. (50) into Eq. (48) and collecting terms oscillating at the same Bogoliubov frequency $\omega_{b,\bm{k}}$ we obtain an eigenvalue equation for $\omega_{b,\bm{k}}$ and the corresponding amplitudes:
$(\eta\mathcal{B}-vp_{x})\begin{pmatrix}U_{b,\bm{k}}\\ V_{b,\bm{k}}\end{pmatrix}=\hbar\omega_{b,\bm{k}}\begin{pmatrix}U_{b,\bm{k}}\\ V_{b,\bm{k}}\end{pmatrix}\,$, (51)
@@
These expressions formally coincide with those derived for the stationary stripe phase [78, 60], except for the additional Doppler term $vp_{x}$ on the left hand side of Eq. (51). As in the $v=0$ case, we seek solutions in the form of Bloch waves:
$\begin{pmatrix}U_{b,\bm{k}}(\bm{r}^{\prime})\\ V_{b,\bm{k}}(\bm{r}^{\prime})\end{pmatrix}=\frac{\mathrm{e}^{\mathrm{i}(\bm{k}+\eta\bm{k}_{c})\cdot\bm{r}^{\prime}}}{\sqrt{V}}\sum_{\text{$\bar{m}$ odd}}\begin{pmatrix}\tilde{U}_{b,\bm{k},\bar{m}}\\ \tilde{V}_{b,\bm{k},\bar{m}}\end{pmatrix}\mathrm{e}^{\mathrm{i}\bar{m}k_{1}x^{\prime}}\,.$ (53)
Here, $\bm{k}$ is the excitation quasimomentum, while the $\tilde{U}_{b,\bm{k},\bar{m}}$’s and $\tilde{V}_{b,\bm{k},\bar{m}}$’s are two-component expansion coefficients. Combining Eqs. (51) and (53) we obtain an infinite-dimensional eigenvalue problem for these coefficients and the corresponding Bogoliubov frequencies $\omega_{b,\bm{k}}$. A detailed discussion of the case $v=0$ can be found in Ref. [56]. At fixed $\bm{k}$, one obtains an infinite set of eigenfrequencies that form a band structure as $\bm{k}$ varies [78, 60]. This justifies the use of an additional subscript, the band index $b$, to label the amplitudes and the frequency of each Bogoliubov mode. The most general solution to Eq. (48) can then be expressed as a sum over all physically distinct excitation bands and quasimomenta.
@@
    complex frequencies occur in complex conjugate pairs. This follows from the fact that the operator $\eta\mathcal{B}-vp_{x}$ differs from its Hermitian adjoint by a unitary transformation: $\eta(\eta\mathcal{B}-vp_{x})\eta^{-1}=(\eta\mathcal{B}-vp_{x})^{\dagger}$ [87];
@@
    each mode has a conjugate counterpart that describes the same physical oscillation. Specifically, if the pair $(U_{b,\bm{k}},V_{b,\bm{k}})$ solves Eq. (51) with frequency $\omega_{b,\bm{k}}$, then the pair $(V_{b,\bm{k}}^{*},U_{b,\bm{k}}^{*})$ is a solution with frequency $\omega_{b,\bm{k}}^{*}$ [87]. However, because of the structure of the Ansatz (50), these two solutions describe the same physical excitation;
@@
[…]
which has two key implications. First, it defines orthogonality between modes belonging to different bands ($b^{\prime}\neq b$) or with different quasimomenta ($\bm{k}^{\prime}\neq\bm{k}$). Second, in the case $b^{\prime}=b$ and $\bm{k}^{\prime}=\bm{k}$, it provides the norm $\mathcal{N}_{b,\bm{k}}$ of a given Bogoliubov mode. Importantly, each physical oscillation corresponds to a pair of solutions with opposite norms. In plotting the Bogoliubov spectrum, such as in Fig. 5, it is customary to include only the frequencies of the positive norm solutions. Furthermore, the system is said to be energetically stable if all positive norm solutions have positive frequency. In the next section, we will show that moving stripe patterns can lose energetic or dynamical stability at sufficiently large dressed spin imbalance. Finally, due to the periodicity of the spectrum, it suffices to consider values of the $x$ component of $\bm{k}$ within the first Brillouin zone, that is, $k_{1}\leq k_{x}<k_{1}$. [^4] Here we depart from the conventions of most previous literature [78, 59, 56, 60, 82, 66], which takes the first Brillouin zone for the excitation quasimomentum as the interval $0\leq k_{x}<2k_{1}$. In contrast, the $y$ and $z$ components are unconstrained.
@@ V.2 Bogoliubov spectrum and sound velocities
To numerically compute the Bogoliubov spectrum, we truncate the expansions (53) by retaining only a finite number of terms. Specifically, we include components with $9\leq\bar{m}\leq 9$, following the approach used in Sec. III.3 for evaluating $\Psi_{0}$. This truncation reduces Eq. (51) to a finite-dimensional eigenvalue problem, which can be readily solved. A detailed discussion of the computational procedure is given in Ref. [56]. In Fig. 5, we display the lowest-lying bands of the computed spectrum for excitation quasimomentum $\bm{k}$ aligned along the $x$ axis. The figure consists of two rows of panels, corresponding to the two values of the Raman coupling considered in this work. Within each row, we vary the value of $s_{p}$ to illustrate four representative cases:
Figure 5: Excitation spectrum of a spin-orbit-coupled BEC in the stripe phase for various values of $s_{p}$: (a1)-(a2) $s_{p}=0$ (stationary stripes), (b1)-(b2) $s_{p}=0.5$, (c1) $s_{p}=0.95$, (c2) $s_{p}=0.75$, and (d1)-(d2) $s_{p}=0.99$. The top-row panels correspond to $\hbar\Omega_{R}/E_{R}=1.0$, while those in the bottom row correspond to $\hbar\Omega_{R}/E_{R}=2.6$. The excitation quasimomentum is taken along the $x$ direction. Only the lowest-lying excitation bands are shown; the band index $b$ is indicated next to each curve. In (c1), the inset shows a magnified view in the vicinity of $k_{x}=0$ on the positive-$k_{x}$ side, where the lowest gapless band exhibits a negative frequency. In all panels, we plot only the real part of the frequency, $\omega_{b,k_{x}}^{\mathrm{R}}$; the imaginary part, $\omega_{b,k_{x}}^{\mathrm{I}}$, is zero except for the lowest gapless band in (c2) (see inset) and in (d2), where it is nonzero but very small and omitted for clarity. The red dashed line in (c2) displays the solution $\omega_{1,k_{x}}^{*}$ in the dynamically unstable region and $\omega_{1,-k_{x}}$ in the dynamically stable region, both obtained from Eq. (51). In panels (d1)-(d2), the black dashed curves represent the two branches of the excitation spectrum in the plane-wave phase, computed using the same parameters as in the case of the stripe-phase spectrum, and with an added Doppler shift (see text for details). The interaction parameters are as in previous figures: $G_{dd}/E_{R}=1.2$ and $G_{ss}/E_{R}=0.32$.
@@
    Spectrum of stationary stripes ($s_{p}=0$), previously computed in Refs. [78, 60]. The two lowest bands are gapless and exhibit linear dispersion near $k_{x}=0$, i.e., the center of the first Brillouin zone. The two zero-frequency modes at $k_{x}=0$ correspond to the Goldstone modes arising from the spontaneous breaking of global-phase and translation symmetries. Recall that, in the regime $g_{dd}>g_{ss}$ considered here, the upper (lower) phonon branch has predominantly density (spin) character [78], corresponding to the superfluid (crystalline) nature of the stripe phase [60]. The roles are interchanged when $g_{ss}>g_{dd}$.
@@
    Spectrum of moving stripes at intermediate dressed spin imbalance ($s_{p}=0.5$). Most features remain similar to the stationary case. However, the bands no longer exhibit reflection symmetry with respect to the center of the Brillouin zone. This is due to the fact that the transformation $k_{x}\to-k_{x}$ transforms the spectrum of a stripe pattern to the one with opposite $s_{p}$. Additionally, the gaps between bands are narrower for moving stripes, and this trend becomes more pronounced as $|s_{p}|$ increases.
@@
    Spectrum in the instability regime ($s_{p}=0.95$ for the top row, $s_{p}=0.75$ for the bottom row). The behavior depends on the Raman coupling. At lower Raman coupling (top row), a portion of the lowest gapless band dips into negative frequency, particularly right of $k_{x}=0$ (or left of $k_{x}=0$ for negative $s_{p}$), as shown in the inset of panel (c1). This signals the onset of energetic instability for moving stripes at large dressed spin imbalance. Physically, this means that in these configurations the increase in energy cost associated with the large bare spin imbalance [see Fig. 3(d1)] is no longer sufficiently compensated by the energy decrease resulting from the reduction of the stripe contrast [Fig. 3(a1)]. Conversely, at higher Raman coupling (bottom row), the system first exhibits energetic instability (now left of $k_{x}=0$), followed by the emergence of a nonzero imaginary part in the lowest band near $k_{x}=0$, as seen in panel (c2). According to the discussion in the previous section, this indicates that moving stripe patterns become dynamically unstable at high Raman coupling and high dressed spin imbalance. We speculate that this dynamical instability would lead to a spatial separation of the two dressed spin components of the stripe order parameter, resulting in the formation of multiple plane-wave domains.
@@
    Spectrum near the melting limit ($s_{p}=0.99$). Unlike the previous panels, which focus on the first Brillouin zone, here we plot the spectrum over the range $4k_{1}\leq k_{x}<2k_{1}$, whose length is three times that of the first Brillouin zone. One observes a significant narrowing of the band gaps, heralding their complete closure at the stripe melting point. At this point, segments from different bands and Brillouin zones merge to form the dispersion relations $\omega_{\pm,\bm{k}}-vk_{x}$ (dashed lines), where $\omega_{\pm,\bm{k}}$ denote the lower (-) and upper (++) branches of the excitation spectrum in the plane-wave phase [69]. [^5] Figure 5(d1)-(d2) may suggest that the Bogoliubov frequencies remain periodic in $k_{x}$ even in the plane-wave phase. However, one must recall that $\bm{k}$ represents the excitation quasimomentum. The spectrum loses periodicity when plotted as a function of the excitation momentum, which is well defined only for $|s_{p}|=1$; see the related discussion in Ref. [60]. The Doppler shift $vk_{x}$ appears because the spectrum is evaluated in the comoving frame. Notably, the density phonon modes (upper gapless branch) of the stripe phase smoothly evolve into the single phonon branch of the plane-wave phase. More intriguingly, the spin phonon modes (lower gapless branch) transform into the well-known roton minimum of the plane-wave excitation spectrum near $k_{x}=-2k_{1}^{\mathrm{PW}}$ (or $k_{x}=+2k_{1}^{\mathrm{PW}}$ for the plane-wave state with negative condensation momentum, corresponding to $s_{p}=-1$[88, 69, 62, 89, 90].